# Memorization to Generalization: Emergence of Diffusion Models from Associative Memory

### v2

@@ Appendix B Experimental Details of the 2D Toy Example
To visualize the energy landscape, we observe the energy at $t=0.15$ which corresponds to $\beta=3.3$.
Sample Clustering. To cluster the generated samples, we use the AgglomerativeClustering algorithm from scikit-learn, which is a part of SciPy [75].
Appendix C Additional Details on the Likelihood of Diffusion Models
For concreteness, we further explain the likelihood computation in Equation 23 within this section. Specifically, we want to highlight the formulation done by [25, 72] which resulted in Equation 23.
Probability Flow ODE. Suppose we have the following forward process
@@ Appendix C Additional Details on the Likelihood of Diffusion Models
which corresponds to Fokker-Planck equation [76]:
@@
where $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)={\mathbf{f}}({\mathbf{x}}_{t},t)-\frac{1}{2}\nabla\cdot\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}-\frac{1}{2}\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$. With careful inspection of Equation 27, it is equal to the Liouville equation if the diffusion term $\tilde{\mathbf{g}}({\mathbf{x}},t)=0$ and essentially, it is the probability flow ODE where
$\begin{split}\mathrm{d}{\mathbf{x}}_{t}&=\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t+\tilde{\mathbf{g}}({\mathbf{x}}_{t},t)\mathrm{d}{\mathbf{w}}_{t}\\ &=\bigg{\{}{\mathbf{f}}({\mathbf{x}}_{t},t)-\frac{1}{2}\nabla\cdot\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}-\frac{1}{2}\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg{\}}\mathrm{d}t\end{split}$ (28)
Using Equation 28, we can derive an appropriate probability flow ODE from the forward process (1) introduced in the main text. For example, we have the following equation
@@
Log-likelihood. Furthermore, if we first take the logarithm of Equation 27 where
@@
where $\nabla_{{\mathbf{x}}}\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ denotes the Jacobian of $\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ and the random variable $\epsilon$ satisfies $\mathbb{E}_{p(\epsilon)}[\epsilon]=0$ and $\text{Cov}_{p(\epsilon)}[\epsilon]=\mathbf{I}$.
@@
However, for the toy model in Section 3, due to its low dimensionality we find that it is possible to not utilize the Hutchinson trace estimator at all, and instead opt to compute the trace of the Jacobian without using the noise estimators — to get the actual trace of the Hessian of the energy.
Appendix D Experimental Details on the Transition Mapping
Dataset Point A Point B Total
(Data Size) (Data Size) (Data Size)
MNIST 4,000 32,000 60,000
CIFAR10 2,000 16,000 50,000
LSUN-CHURCH 2,000 16,000 126,227
FASHION-MNIST 4,000 16,000 60,000
Table 1: Table showing the critical points $A$ and $B$ of the memorization-generalization transition for each dataset. We use 30 linearly spaced points between $A$ and $B$ to finely characterize the transition plots shown in Figure 5.
@@ Appendix D Experimental Details on the Transition Mapping
Points Selection. For the computation of Figure 5, we follow the experiment setup of [45] and conduct a sparse search starting at data size $500$ and doubling it all the way to the total data size $|\mathsf{S}|$, i.e., $K=500,1000,2000,\dots,|\mathsf{S}|$. We then identified two transitional critical points, $A$ and $B$, see Table 1, to conduct a more thorough search to elucidate the memorization-generalization transition. Point $A$ indicates the initial drop in memorization while $B$ signals the plateauing of memorization. To capture the finer details of this transition, we performed a search of $30$ inclusive linearly spaced points, from point $A$ to point $B$. For regions outside of the transition, using linear spacing, we sample $5$ points from $|\mathsf{S}|=2$ to point $A$, and another $5$ points from $B$ to the total dataset size $|\mathsf{S}|$, inclusively. In other words, we train a separate model for each selected data size and generate the corresponding evaluation and synthetic sets — keeping the same configuration for all models, see Table 3. Then, we computed the memorization, spurious, and generalization fractions using equations (13), (14), and (15). The selection process for the values $\delta_{m}$ and $\delta_{s}$ is explained below.
Dataset $\delta_{m}$ $\delta_{s}$ $d(\cdot)$
CIFAR10 0.03 0.018 LPIPS
LSUN-CHURCH 0.12 0.05 LPIPS
MNIST 4.5 3.5 $L_{2}-$distance
FASHION-MNIST 4.5 2.5 $L_{2}-$distance
Table 2: Table displaying memorized and spurious threshold values in the computation of the transition plots in Figure 5.
@@
Selection of the Distance Metric. For high-dimensional datasets, CIFAR10 [63] and LSUN-CHURCH [66], we utilized LPIPS [77] with the AlexNet [78] backbone, as the function $d(\cdot,\cdot)$ for both memorization (13) and spurious (14) detection metrics. We selected this approach since it is a commonly used perceptual metric that compares the similarity between two images based on their feature representations. Moreover, it has been shown to better align with human judgment of visual similarity, making it ideal for assessing the quality and diversity of generated samples in these high-dimensional image datasets [77]. For simpler datasets like MNIST [64] and FASHION-MNIST [65], where images are single-channel and less complex, we found that $L_{2}$-distance suffices for both memorization and spurious detection metrics.
@@
Selection of Threshold Values $\delta_{m}$ and $\delta_{s}$. The detection thresholds, $\delta_{m}$ and $\delta_{s}$, were set based on the chosen distance metric $d(\cdot,\cdot)$ and visual inspection of the distance histograms for bimodality as mentioned in Section 4 of the main text. Moreover, these thresholds were chosen to reflect the varying visual complexity and feature richness of the datasets, and their values for each dataset are shown in Table 2. For CIFAR10, we utilized the histograms obtained at $K=7310$, which demonstrate strong bimodality shape, to tune our threshold values to minimize the number of spurious samples classified as memorized samples and vice versa, via manual inspection. For manual inspection, a general rule we followed — it is better to have more false positives in the generalized set than the memorized set since spurious samples can be seen as early-generalized samples. Thus, we observe the least-$5$ spurious samples and ensure that they do not resemble memorized samples. We performed the same process for memorized samples. The entire process is then repeated for the other three datasets. We utilized $K=4896$ for LSUN-CHURCH, $K=21379$ for MNIST, and $K=7724$ for FASHION-MNIST to select $\delta_{m}$ and $\delta_{s}$. For more details on the selected histograms, please refer to Figs. (4), (14), (12), and (13).
@@
Dataset Initial Latent Channel Multipliers Number of Parameters Batch Size Training Iterations
@@
LSUN-CHURCH 128 (1, 1, 2, 2, 4, 4) 109.7M 64 800,000
FASHION-MNIST 128 (1, 2, 2) 24.5M 128 400,000
Table 3: Table displaying both model and training configurations for each dataset.
Model and Training Details. For each point in our transition plots in Figure 5, we train a DDPM-based diffusion model, where the score model is a PixelCNN++ based U-Net [79, 80]. We keep the variances, $\beta_{\text{min}}=10^{-4}$ and $\beta_{\text{max}}$ = $2\times 10^{-2}$, timesteps $T=1000$, and learning rate $lr=$ $2\times 10^{-4}$ for all models and datasets. Each model has $2$ residual blocks [81] for each down- and up- sampling layer, while an attention block is placed at 16×16$\times$16 × resolution. We only modified the channel multipliers for each model based on the complexity of the dataset, see Table 3. If the dataset size is smaller than the specified batch size, we take the batch size to be equal to that small dataset size. For generation or inference, we use the exponential moving average (EMA) of each trained model, as delineated in [23], which was obtained with the decay value set as $0.9999$ during training. Please note that we did not use random flipping in the training of our models, since we want our measurements to reflect the training data size at best as random flipping implicitly increases the number of patterns that the models see during training. However, we did use dropout (of value $0.1$) for the training of CIFAR10, MNIST, and FASHION-MNIST models. Lastly, for each of the training set $\mathsf{S}_{\alpha}$, where $\alpha=1,\dots,M$, they are split from the original dataset given a specific size, using the same random seed value of $3407$.
Figure 9: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for CIFAR10 [63] and various model’s widths. As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row). Additionally, with a smaller number of parameters, the memorization-generalization transition happens at an earlier stage (or smaller $K$), since the model has a memorization capacity. In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data. The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table 3.
Additional Experiments. Inspired by [45, 46], we extend our investigation on the memorization-generalization transition by studying the effects of varying the parameter size of the diffusion model for CIFAR10 [63] and LSUN-CHURCH [66]. The results of this experiment are recorded in Figs. (9) and (10), respectively. Specifically, we trained two additional sets of 38 diffusion models, with the initial latent dimensions of 64 and 96 respectively, and repeat the same experimental setup delineated above, see also Table 3.
For CIFAR10 and the U-Net (64) models, we identified the critical points, $A$ and $B$, to be $500$ and $8000$ respectively. Meanwhile, for the U-Net (96) models, these critical points are $2000$ and $16000$. To select $\delta_{m}$ and $\delta_{s}$, we followed the same process detailed above by observing the distance histograms to see where bimodality occurs with respect to the distance value, and manually inspecting the images to ensure there is a minimal amount of false positives in memorized and spurious samples. For U-Net (64) models, the threshold values $\delta_{m}$ and $\delta_{s}$ are both selected to be $0.02$ based on the bimodality of the distance histograms observed at $K=2568$ — meanwhile, for U-Net (96), the threshold values $\delta_{m}$ and $\delta_{s}$ are chosen as $0.03$ and $0.02$ according to distance histograms observed at $K=3931$.
Similarly, for the LSUN-CHURCH models, we identified the critical points, $A$ and $B$, to be $1000$ and $8000$ for the U-Net (64) models; and for the larger U-Net (128) models, $4000$ and $16000$ as the critical points. The threshold values, $\delta_{m}$ and $\delta_{s}$, are both selected to be $0.04$ after observing the bimodality of the distances at $K=3413$ — meanwhile, for U-Net (128), these values are selected to be $\delta_{s}=0.04$ and $\delta_{m}=0.1$ after observing the distance histograms at $K=5241$.
Figure 10: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for LSUN-CHURCH [66] and various model’s widths. As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row). Additionally, with a smaller number of parameters, the memorization-generalization transition happens at an earlier stage (or smaller $K$), since the model has a memorization capacity. In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data. The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table 3.
With the smaller U-Net (64) models, we observe the memorization-generalization transition happens at an earlier stage with respect to the larger U-Net (128) models, where the peak of spurious samples appears at $K=2568$ instead of $K=5862$ for CIFAR10 and $K=1000$ instead of $K=4000$ for LSUN-CHURCH. Similarly, for CIFAR10 and its U-Net (96) models, the peak of spurious samples happens at an early stage of $K=3931$ instead. Overall, these results indicate that as the parameter size of the diffusion model increases, the memorization capacity of the model is increased and thus, generalization is delayed since the model can memorize more of its training data denoted by the appearance of spurious patterns happening at a larger training data size $K$. In contrast, when the memorization capacity is reduced, generalization is initiated at an earlier stage (or smaller training data size) with respect to when the parameter size of the model is larger. Most importantly, the emergence of spurious patterns remains consistent despite the variations in the parameter size of the diffusion model.
Figure 11: Memorization-Generalization transition plots for the CIFAR10 dataset. The left figure depicts the memorization-generalization transition computed with 4×4$\times$4 × synthetic sets for various training data sizes. Meanwhile, right figure shows the transition computed using 8×8$\times$8 × synthetic sets for the same training data sizes. With the larger synthetic sets, more spurious patterns are detected at the peak ($K=5862$). However, the overall trend of these patterns quickly rise, during the co-existence of memorization and generalization, and fall as generalization becomes the dominant phase, is maintained. Both figures use the same memorized and spurious threshold values detailed in Table 2.
Additionally, to verify the consistency of the memorization-generalization transition, we explored the influence of larger synthetic sets and recorded our results in Figure 11. Specifically, we repeated our experiment done in Figure 5 for CIFAR10 using 8×8$\times$8 × synthetic sets — which was originally done using 4×4$\times$4 × synthetic sets. Based on Figure 11, we observe a slight increase in the fraction of spurious samples when using the larger 8×8$\times$8 × synthetic sets. Specifically, at $K=5862$, the fraction increases from $0.04$ to roughly $0.06$ with respect to the synthetic set size. Nonetheless, the trend of memorization-generalization, with the emergence of spurious samples, is still maintained regardless of the multiplier for the synthetic size — highlighting the consistency of the spurious phenomenon in diffusion models in relation to the training data size.
Hardware Details. The training of the models detailed in Table 3 was done using NVIDIA Tesla V100 GPUs, where a single GPU was used to train each model. Each GPU has 32GB of memory and is linked with IBM Power9 processors which clocked at 3.15 GHz. Each CIFAR10 model took roughly 18, 20, and 22 hours to train for different initial latent dimensions, e.g., 64, 96, 128, respectively. Meanwhile, each LSUN-CHURCH model requires roughly 48, 56, and 103 hours of training with respect to the model’s initial latent dimension. In contrast, FASHION-MNIST and MNIST models require approximately 17 hours to be fully trained (per model). Please note that these durations are with respect to the required training iterations for each model and their corresponding dataset detailed in Table 3. Please note that we utilized the same hardware for our experiments as well.
Appendix E Additional Visualizations of Different Sample Types Across the Transition
E.1 LSUN-CHURCH

### v3

@@ A.2  Experimental Details
To visualize the energy landscape, we observed the energy at $t=0.15$ which corresponds to $\beta=3.3$.
Model and Training Details. For each training data size $K$, we used an MLP-based network to estimate the score $s_{\theta}({\mathbf{x}}_{t},t)$. The general architecture includes (1) a Fourier random feature (FRF) timestep embedding layer and a linear layer which projects the concatenation of FRF timestep embedding and input into a latent dimension of 256; (2) an encoder which consists of four non-convolutional residual blocks (using the same latent dimension) with Swish [47] activation in between; (3) a decoder which consists of the same number of residual blocks; and (4) a linear block which projects the latent variable back to the 2D space.
The network was trained in continuous time with $\sigma=1$ using the objective function from [22, 62]. Optimization is performed using the Adam [34] optimizer with a learning rate $lr=10^{-4}$. The batch size is set to $\textbf{min}(K,500)$. All models were trained for $800000$ iterations with the maximum batch size of $500$. Meanwhile, we used Euler-Maruyama discretization method [35], with 1000 discretized steps, to solve the reverse SDE (20) to generate samples.
Sample Clustering. To cluster the generated samples, we used the AgglomerativeClustering algorithm from scikit-learn, a part of SciPy [68].
A.3  Log-Likelihood of Diffusion Models
For concreteness, we provide the likelihood computation of Eq. (24) within this section to highlight the formulation of Eq. (24) done by [62, 13].
Probability Flow ODE. Suppose we have the following forward process:
@@ A.3  Log-Likelihood of Diffusion Models
which corresponds to the Fokker-Planck equation [46]:
@@
where $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)={\mathbf{f}}({\mathbf{x}}_{t},t)-\frac{1}{2}\nabla\cdot\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]-\frac{1}{2}\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$. With careful inspection of Eq. (28), it is equal to the Liouville equation if the diffusion term $\tilde{\mathbf{g}}({\mathbf{x}},t)=0$ and essentially, it is the probability flow ODE where
$\begin{split}\mathrm{d}{\mathbf{x}}_{t}&=\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t+\underset{0}{\underbrace{\tilde{\mathbf{g}}({\mathbf{x}}_{t},t)\mathrm{d}{\mathbf{w}}_{t}}}\\ &=\bigg\{{\mathbf{f}}({\mathbf{x}}_{t},t)-\frac{1}{2}\nabla\cdot\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]-\frac{1}{2}\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg\}\mathrm{d}t\end{split}$ (29)
Using the ODE (29), we can derive an appropriate probability flow ODE from the forward process detailed in the main text. For example, we have the following equation
@@
Log-likelihood. Furthermore, if we first take the logarithm of Eq. (28) where
@@
where $\nabla_{{\mathbf{x}}}\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ denotes the Jacobian of $\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ and the random variable $\epsilon$ satisfies $\mathbb{E}_{p(\epsilon)}[\epsilon]=0$ and $\text{Cov}_{p(\epsilon)}[\epsilon]=I$.
@@
However, for the toy example, due to its low dimensionality we found that it is possible to not utilize the Hutchinson trace estimator at all, and instead opted to compute the trace of the Jacobian without using the noise estimators.
Appendix B Memorization-to-Generalization Transition
B.1  Selection of Points
For the computation of fig.˜4 in the main text, we followed the experiment setup of [74] and conducted an initial sparse search of the transition, starting at data size $500$ and doubling it all the way to the total data size $|\mathsf{S}|$, i.e., $K\in\{500,1000,2000,\dots,|\mathsf{S}|\}$. We then identified two transitional critical points, $A$ and $B$, see Table (1), to conduct a more thorough search to elucidate the memorization-to-generalization transition.
@@ B.1  Selection of Points
Here, point $A$ indicates the initial drop in memorization while $B$ signals the plateauing of memorization. To capture the finer details of this transition, we performed a search of $30$ inclusive linearly spaced points, from point $A$ to point $B$. For regions outside of the transition, using linear spacing, we sampled $5$ points from $|\mathsf{S}|=2$ to point $A$, and another $5$ points from $B$ to the total dataset size $|\mathsf{S}|$, inclusively. In other words, we trained a separate model for each selected data size and generated the corresponding evaluation and synthetic sets while keeping the same configuration for all models, see Table (3). Then, we computed the memorization, spurious, and generalization fractions using the detection metrics in Sec. (4). The selection process for the values $\delta_{m}$ and $\delta_{s}$ is explained below.
Table 1: A table showing the critical points $A$ and $B$ of the memorization-generalization transition for each dataset.11footnotetext: Note: We use 30 linearly spaced points between $A$ and $B$ to finely characterize the memorization-to-generalization transition.
Training Data Size
Dataset Point AA Point $B$ Total
MNIST 4000 32000 60000
CIFAR10 2000 16000 50000
LSUN-CHURCH 2000 16000 126227
FASHION-MNIST 4000 16000 60000
B.2  Selection of Distance Metric
@@
For high-dimensional datasets, CIFAR10 [37] and LSUN-CHURCH [75], we utilized LPIPS [76] with the AlexNet [38] backbone, as the function $d(\cdot,\cdot)$ for both memorization and spurious detection metrics. We selected this approach since it is a commonly used perceptual metric that compares the similarity between two images based on their feature representations. Moreover, it has been shown to better align with human judgment of visual similarity, making it ideal for assessing the quality and diversity of generated samples in these high-dimensional image datasets [76]. For simpler datasets like MNIST [17] and FASHION-MNIST [73], where images are single-channel and less complex, we found that $L_{2}$-distance suffices for both memorization and spurious detection metrics.
B.3  Selection of Threshold Values
@@
The detection thresholds, $\delta_{m}$ and $\delta_{s}$, were set based on the chosen distance metric $d(\cdot,\cdot)$ and visual inspection of the distance histograms for bimodality as mentioned in Sec. (4). Moreover, these thresholds were chosen to reflect the varying visual complexity and feature richness of the datasets, and their values for each dataset are shown in Table (2). Specifically, for each training data size $K$ of each dataset, we manually inspected the bimodality of the olive and gray histograms to select $\delta_{m}$ and $\delta_{s}$. Moreover, to better tune these thresholds at each $K$, we manually inspected the least $5$ spurious samples, ensuring that they do not resemble memorized samples, while doing the same for memorized samples. If the resemblance of one set to another is too strong, we decreased the corresponding threshold. Please refer to Figs. (10)-(12) below and fig.˜3 in the main text for examples of the histograms. Please note that the bimodality of the gray histograms in FASHION-MNIST might not be significant, but it is enough to select their respective $\delta_{s}$.
B.4  Model and Training Details
For each point in our transition plots of fig.˜4, we trained a DDPM-based diffusion model, where the score model is a PixelCNN++ based U-Net [64, 53]. We kept the variances, $\beta_{\text{min}}=10^{-4}$ and $\beta_{\text{max}}$ = $2\times 10^{-2}$, timesteps $T=1000$, and learning rate $lr=$ $2\times 10^{-4}$ for all models and datasets. Each model has $2$ residual blocks [21] for each down-sampling and up-sampling layer, while an attention block is placed at $16\times$ resolution. We only modified the channel multipliers for each model based on the complexity of the dataset, see Table (3). If the training data size $K$ is smaller than the specified batch size, the batch size is set to be equal to $K$. For generation or inference, we used the exponential moving average of each trained model, as delineated in [22], which is obtained with the decay value set as $0.9999$ during training. Please note that we did not use random flipping in the training of our models, since we want our measurements to reflect the training data size at best as random flipping implicitly increases the number of patterns that the models see during training. However, we did use dropout (of value $0.1$) for the training of CIFAR10, MNIST, and FASHION-MNIST models. For LSUN-CHURCH dataset, the images were center-cropped and down-sampled to $64\times 64$ resolution. Lastly, for each of the training set $\mathsf{S}_{\alpha}$, where $\alpha=1,\dots,M$, they are split from the original dataset given a specific size, using the same random seed value of $3407$.
B.5  Additional Results
Figure 7: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for CIFAR10 [37] and various model’s widths. As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row). Additionally, with a smaller number of parameters, the memorization-generalization transition happens at an earlier stage (or smaller $K$), since the model has a memorization capacity. In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data. The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table (3).
Inspired by [74, 20], we extend our investigation on the memorization-generalization transition by studying the effects of varying the parameter size of the diffusion model for CIFAR10 [37] and LSUN-CHURCH [75]. The results of this experiment are recorded in Figs. (7)-(8) for these two datasets respectively. Specifically, we trained two additional sets of 38 diffusion models, with the initial latent dimensions of 64 and 96 respectively, and repeated the same experimental setup delineated above, see also Table (3).
For CIFAR10 and the U-Net (64) models, we identified the critical points, $A$ and $B$, to be $500$ and $8000$ respectively. Meanwhile, for the U-Net (96) models, these critical points are $2000$ and $16000$. To select $\delta_{m}$ and $\delta_{s}$, we followed the same process detailed above by observing the distance histograms to see where bimodality occurs with respect to the distance value, and manually inspecting the images to ensure there is a minimal amount of false positives in memorized and spurious samples. Similarly, for the LSUN-CHURCH models, we identified the critical points, $A$ and $B$, to be $1000$ and $8000$ for the U-Net (64) models; and for the larger U-Net (128) models, $4000$ and $16000$ as the critical points. Please refer to Table (2) for the threshold values.
Figure 8: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for LSUN-CHURCH [75] and various model’s widths. As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row). Additionally, with a smaller number of parameters, the memorization-generalization transition happens at an earlier stage (or smaller $K$), since the model has a memorization capacity. In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data. The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table (3).
With the smaller U-Net (64) models, we observe the memorization-generalization transition happens at an earlier stage with respect to the larger U-Net (128) models, where the peak of spurious samples appears at $K=2568$ instead of $K=5862$ for CIFAR10, and $K=4862$ instead of $K=6086$ for LSUN-CHURCH. Overall, these results indicate that as the parameter size of the diffusion model increases, the memorization capacity of the model is increased and thus, generalization is delayed since the model can memorize more of its training data denoted by the appearance of spurious patterns happening at a larger training data size $K$. In contrast, when the memorization capacity is reduced, generalization is initiated at an earlier stage (or smaller training data size) given a larger parameterized model. Most importantly, the emergence of spurious patterns remains consistent despite the variations in the parameter size of the diffusion model: these patterns quickly rise and fall at the boundaries of the memorization-to-generalization transition.
Figure 9: Memorization-Generalization transition plots for the CIFAR10 dataset. The column (a) figure depicts the memorization-generalization transition computed with $4\times$ synthetic sets for various training data sizes. Meanwhile, the column (b) figure shows the transition computed using $8\times$ synthetic sets for the same training data sizes. With the larger synthetic sets, more spurious patterns are detected around the peak of the transition. However, the overall trend of these patterns quickly rise, during the co-existence of memorization and generalization, and fall as generalization becomes the dominant phase, is maintained. Both figures use the same memorized and spurious threshold values detailed in Table (2).
Additionally, to verify the consistency of the memorization-generalization transition, we explored the influence of larger synthetic sets and recorded our results in Fig. (9). Specifically, we repeated our experiment done in Fig. (4) for CIFAR10 using $8\times$ synthetic sets, which was originally done using $4\times$ synthetic sets. Based on these results, we observe a slight increase in the fraction of spurious samples when using the larger $8\times$ synthetic sets. Specifically, at $K=5862$, the fraction increases from $0.04$ to roughly $0.06$ with respect to the synthetic set size. Nonetheless, the trend of memorization-generalization, with the emergence of spurious samples, is still maintained regardless of the multiplier for the synthetic size. This aspect highlights the consistency of the spurious phenomenon in diffusion models in relation to the varying of training data size.
Table 2: Table displaying the range of the memorized and spurious threshold values in the computation of the transition plots in Fig. (4) of the main text.11footnotetext: The full set of threshold values is provided alongside with the code.
Range [min, max]
Dataset Initial Latent Dim. $\delta_{m}$ $\delta_{s}$
MNIST 128 $[2,4.5]$ $[2,4]$
FASHION-MNIST 128 $[2,2.5]$ $[2.5,3]$
64 $[0.01,0.04]$ $[0.015,0.025]$
CIFAR10 96 $[0.015,0.04]$ $[0.015,0.025]$
128 $[0.015,0.035]$ $[0.015,0.025]$
64 $[0.06,0.15]$ $[0.01,0.1]$
LSUN-CHURCH 96 $[0.1,0.125]$ $[0.05,0.06]$
128 $[0.1,0.12]$ $[0.04,0.06]$
Table 3: A table displaying both model and training configurations for each dataset.
Configurations
@@ B.5  Additional Results
Dataset Initial Latent Dim. Channel Multipliers Num. of Param. Batch Size Training Iterations
@@
128 (1, 1, 2, 2, 4, 4) 109.7M 64 800,000
B.6  Additional Visualizations of Distance Histogram